OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, find the limit. Equation in comments

OpenStudy (anonymous):

\[a_n= \frac{ (-1)^n }{ 5\sqrt{n} }\] \[\lim_{n \rightarrow \infty} a_n = ? \]

OpenStudy (anonymous):

What do you think? ^.^

OpenStudy (anonymous):

I think it converges but i don't know how to do it

OpenStudy (anonymous):

Yep, it definitely converges :)

OpenStudy (anonymous):

how do i figure out where it converges?

OpenStudy (anonymous):

Use a little something called... the Squeeze Theorem ^.^

OpenStudy (anonymous):

Maybe you'll understand better when you take a look at this... \[\Huge \frac{-1}{5\sqrt n}\le\frac{(-1)^n}{5\sqrt n}\le\frac1{5\sqrt n}\]

OpenStudy (anonymous):

And this is in fact, true, for all \(\large n \in \mathbb{N}\)

OpenStudy (anonymous):

so it would be 1/5?

OpenStudy (anonymous):

No. Obviously the sequences on the left and right are both convergent, right? But to where?

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{ 1 }{ 5\sqrt{n} }\]

OpenStudy (anonymous):

Yeah... so? What's the limit?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

That's right :) So, whatever the limit of your sequence is, it can't be greater than 0, right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Now, what's the limit of the left sequence?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

That's right :) So, since it was the left sequence, your original sequence cannot have a limit that is LESS an zero, right?

OpenStudy (anonymous):

right. So it converges at 0

OpenStudy (anonymous):

Yes. Because its limit cannot be greater than zero, and also cannot be less than zero... It got SQUEEZED by two convergent sequences, so to speak ^.^

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

NO problem ^.^